OpenStudy (anonymous):

23. i2 = √(-1) True False

jimthompson5910 (jim_thompson5910):

i = sqrt(-1) i^2 = (sqrt(-1))^2 i^2 = -1

OpenStudy (anonymous):

so is it true?

jimthompson5910 (jim_thompson5910):

i^2 = -1, is that equal to sqrt(-1) ?

OpenStudy (anonymous):

oh no! its not

jimthompson5910 (jim_thompson5910):

so it's false

OpenStudy (anonymous):

can you hellp me with one more? The minimum value is the smallest y-value of a function. true or false?

jimthompson5910 (jim_thompson5910):

what do you think?

OpenStudy (anonymous):

true

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

okay and one more Complex numbers can be graphed on the xy coordinate plane? is this false?

OpenStudy (anonymous):

i mean true?*

jimthompson5910 (jim_thompson5910):

it's true

OpenStudy (anonymous):

woo thank you! for your help!

jimthompson5910 (jim_thompson5910):

np